# Boaz Moerman

### v2

@@ 1.4. Evidence for the conjecture
In an upcoming work [Moe25toric], we prove Conjecture 1.4 for split toric varieties, providing the first counting results on $\mathcal{M}$-points in general.
Theorem 1.12.
Let $(X,M)$ be a smooth proper pair over $\mathbb{Q}$, where $X$ is a smooth proper split toric variety and the divisors $D_{1},\dots,D_{n}$ are torus-invariant. Let $\mathcal{X}$ be the $\mathbb{Z}$-integral model of $X$ induced by the fan of $X$ and let $(\mathcal{X},\mathcal{M})$ be the corresponding integral model of $(X,M)$. Then Conjecture 1.4 is true for $(X,M)$ for any big and nef divisor class $L$ on $X$ with toric metrization $\mathcal{L}$ (as defined in [BaTs96, Theorem 2.1.6]).
In Section 9, we give further evidence for the conjecture by showing that the conjecture is compatible with the the various results on $\mathcal{M}$-points of bounded height [Str21, ShSt24, Ito25, Ara25]. In Theorem 9.1 we give an asymptotic formula for the number of elements with $m$-full norm in any Galois extension of $\mathbb{Q}$, under assumption of Conjecture 1.4. Furthermore, we show that this asymptotic agrees with the result by Streeter [Str21, Theorem 1.1], whenever the hypotheses of his result are satisfied. In order to compare our leading constant with the one found by Streeter, we prove Lemma 7.8, giving a description of the algebraic Brauer group of pairs $(X,M)$ over toric varieties in terms of automorphic characters.
@@ 1.5. Structure of the paper
In Section 3, we introduce pairs and $\mathcal{M}$-points, as well as various special cases of these such as Campana points. In Section 4.2, we introduce the Picard group of a pair, as well as its effective cone and its canonical divisor. In the same section, we also prove several results to aid in the computation of the Fujita invariant and the $b$-invariant. In Section 5 we define when a pair is rationally connected, and motivate its role in Conjecture 1.4. Section 6 is devoted to the study of the log-canonical divisor and quasi-Campana pairs, which are pairs loosely resembling the pairs for Campana points. For such pairs, we compute the Fujita-invariant and the $b$-invariant, yielding Theorem 1.7 as a special case. In Section 7, we introduce the fundamental group and the Brauer group of a pair, and we study the latter for pairs over toric varieties. In Section 8, we define the leading constant in Conjecture 1.4 using the Brauer group from the previous section. Finally, in Section 9, we prove that the previous results known on $\mathcal{M}$-points of bounded height are compatible with Conjecture 1.4.
@@ Acknowledgments
This article is based on my PhD thesis [Moe24], which I wrote at Utrecht University under supervision of Marta Pieropan. I would like to thank her for her unwavering support for my work and her various suggestions. I would also like to thank Alec Shute, Sam Streeter, Tim Santens and Damaris Schindler for the useful discussions we had during the workshop on Campana points on toric varieties in Bristol, as well as on other occasions. Finally, I would like to thank Sam Streeter, Tim Santens and Dan Loughran for discussions on the leading constant.
@@ 3. Pairs and $\mathcal{M}$-points
In this section we introduce pairs and the $\mathcal{M}$-points on such pairs. This notion will be based on the notion of pair previously considered in [Moe24, Thesis]. In this article, we work with a more general notion. The differences are are as follows:
@@
    The parameter set $M$ does not only describe which intersection multiplicities $\mathcal{M}$-points should have, it also can impose to which intersection strata these multiplicities should correspond to.
@@
The original notion of pair corresponds to a split pair which is divisorial, as defined in this article. There are several reasons for considering more general pairs. For instance, on a non-split smooth toric variety, the torus-invariant prime divisors are not smooth in general, but their geometric components are smooth. This is significant, as our generalization of Manin’s conjecture requires the divisors $D_{1},\dots,D_{n}$ defining the pair to be smooth. We start by introducing split pairs, which are pairs for which the boundary divisors are already defined over the base field. These pairs are the pairs considered in [Moe24, Thesis].
3.1. Split pairs
In this section we will introduce split pairs and their $\mathcal{M}$-points. We will introduce these notions in large generality, using the language of scheme theory. While we are mainly interested in $\mathcal{M}$-points over rings of integers, we will also need to consider $\mathcal{M}$-points over $\mathbb{P}^{1}$ in Definition 5.1 to define what it means for a pair to be rationally connected. Furthermore, the general framework presented here has nice functorial properties as seen in Remark 3.18.
In general, we want to consider divisors which need not be not defined over the base, but only over a Galois cover. Let us recall the definition of a Galois cover as in [Stacks, Tag 03SF], generalizing Galois extensions of a field.
@@ 3.1. Split pairs
For the remainder of the section, all schemes will be schemes over $B$, and let $A\rightarrow B$ be a fixed Galois cover with Galois group $G$.
@@
Let $P\colon Y\rightarrow X$ be a morphism of schemes and let $D$ be an effective divisor on $X_{A}$. For any prime (Cartier) divisor $v$ on $Y_{A}$, let $\mathcal{O}_{v}$ be the local ring of $Y_{A}$ along $v$ and let $P_{v}\colon\operatorname{Spec}\mathcal{O}_{v}\rightarrow X$ be the morphism obtained by composing $P$ with the natural morphism $\operatorname{Spec}\mathcal{O}_{v}\rightarrow Y$. We define the (local) intersection multiplicity of $P$ and $D$ at $v$ to be the extended integer $n_{v}(D,P)\in\overline{\mathbb{N}}$ such that $\operatorname{Spec}\mathcal{O}_{v}\times_{X}D=\operatorname{Spec}\mathcal{O}_{v}/(\pi_{v})^{n_{v}(D,P)}$, where $\pi_{v}$ is an uniformizer of $\mathcal{O}_{v}$ and we set $(\pi_{v})^{\infty}=0$. Similarly, for an irreducible component $v$ of $Y_{A}$, we set
@@
Note that if the image of $P$ is contained in $D$, then $n_{v}(D,P)=\infty$ for every Cartier divisor $v$ on $Y$. If $D$ is Cartier, the image of $P$ is not in $D$ and $Y$ is integral, then $P^{-1}D$ is the pullback divisor $P^{*}D$ [Stacks, Tag 02OO] and $n_{v}(D,P)$ is the coefficient of $v$ in $P^{*}D$.
@@
Thus $n_{v}(D,P)$ is the local intersection multiplicity of $C$ and $D$ in a point $v$ as defined in [Har77, Chapter V].
@@
If $X=\mathbb{A}_{\mathbb{Z}}^{2}$, $D_{1}=\{y=0\}$, and $D_{2}=\{y=x^{2}-1\}$, then an integral point $P=(a,b)\in\mathbb{Z}^{2}$ satisfies $n_{p}(D_{1},P)=v_{p}(b)$ and $n_{p}(D_{2},P)=v_{p}(b-a^{2}+1)$ for every prime number $p$, where $v_{p}$ is the $p$-adic valuation. In particular, it follows that $n_{p}(D_{1},P),n_{p}(D_{2},P)>0$ if and only if $b=0\bmod p$ and $a=\pm 1\bmod p$. The intersection multiplicities at $p$ tells us whether the point reduces to $D_{1}\cap D_{2}=\{(1,0),(-1,0)\}$ modulo $p$. However, it does not tell us to which component it reduces to. In order to keep track of this, we introduce the following notation.
@@
We write $\overline{\mathbb{N}}^{n}_{\mathcal{C}}$ for the set of tuples $(\mathbf{m},c)$, where $\mathbf{m}\in\overline{\mathbb{N}}$ and $c\in\mathcal{C}_{\mathbf{m}}$, on which $G$ acts by $\sigma(\mathbf{m},c)=(\sigma(\mathbf{m}),\sigma(c))$.
@@
The strata $c$ in Definition 3.6 correspond to the faces of the corresponding Clemens complex (see e.g. [San23-2]) whenever $X$ is a variety and $\sum_{i=1}^{n}D_{i}$ is a normal crossings divisor.
@@
The previous example has a straightforward generalization to toric varieties, see e.g. [Moe24, §5].
@@
Let us compute $\operatorname{mult}^{\mathcal{C}}_{v}(P)$ in Example 3.5 for integral points $P=(a,b)\in\mathbb{Z}^{2}$. For every prime number $p$ we write $\mathbf{m}_{p}=(v_{p}(b),v_{p}(b-a^{2}+1))$. If at least one of the two valuations is positive, we have
$\operatorname{mult}_{p}^{\mathcal{C}}(a,b)=\{(\mathbf{m}_{p},X)\}.$
If both valuations are positive and $p$ is odd, then we have
@@
where $\overline{a}=a\bmod p\in\{1,-1\}$. On the other hand, if both valuations are positive and $p=2$, then
$\operatorname{mult}_{2}^{\mathcal{C}}(a,b)=\big\{\big((\mathbf{m}_{2},\{(1,0)\}\big),\big(\mathbf{m}_{2},\{(-1,0)\}\big)\big\}.$
@@
If $Y$ is irreducible, then the condition on the generic point is automatically satisfied if the image of $Y$ is not contained in any of the divisors.
@@
The assumption that $Y$ is regular and Noetherian implies that every divisor on $Y$ is a sum of prime Cartier divisors by [Stacks, Tag 0BCP]. Therefore, the values $\operatorname{mult}_{v}(P)$ for all $v$ as above together determine the inverse images $P^{-1}D_{1},\dots,P^{-1}D_{n}\subset Y_{A}$.
In many cases, we do not impose any condition on the components $c$. To make the notation less cumbersome in this case, we introduce the following definition.
@@
The definition of a divisorial split pair given here strongly resembles the definition of a pair previously used by the author in [Moe24, Definition 3.1], but there are two minor differences. In this article, we assume for simplicity that $D_{1},\dots,D_{n}$ are divisors rather than arbitrary closed subschemes. Furthermore, in the other article, we imposed a weak condition on the set $\mathfrak{M}$ ensuring that the $\mathcal{M}$-points over $\mathbb{Z}$ are a subset of the $\mathcal{M}$-points over $\mathbb{Q}$. This condition is not necessary in the framework used in this paper, as the $\mathcal{M}$-point condition used here also takes generic points into account.
@@
If the divisors $D_{1},\dots,D_{n}$ are chosen such that $D_{1}+\dots+D_{n}$ is a strict normal crossings divisor as in [Stacks, Tag 0BI9], then [Stacks, Tag 0BIA] implies that the connected components of the intersections $\bigcap_{i\in I}D_{i}$ are all irreducible for all nonempty $I\subset\{1,\dots,n\}$. Thus it follows that for any scheme $Y$ over $B$ with a prime divisor (or connected component) $v$, the function $\operatorname{mult}_{v}^{\mathcal{C}}$ can be regarded as a function $X(Y)\rightarrow\overline{\mathbb{N}}^{n}_{\mathcal{C}}$. In particular, the condition at $v$ for $P\colon Y\rightarrow X$ to be an $M$-point simplifies to $\operatorname{mult}^{\mathcal{C}}_{v}(P)\in\mathfrak{M}_{\mathcal{C}}$.
@@
for $a_{v}\in\overline{\mathbb{N}}$. Here $a_{v}$ is the multiplicity of $v^{\prime}$ in the pullback of $v$ to $Y^{\prime}_{A}$ if $f(Y^{\prime})\not\subset v$ and $a_{v}=\infty$ otherwise.
@@ 3.2. Pairs over number fields
Let $(X,M)$ be a pair over a number field $K$, where $X$ is a proper variety. For an integral model $\mathcal{X}$ of $X$ over $\mathcal{O}_{S}$, where $S$ is a finite set of places of $K$, there is a natural choice of an $\mathcal{O}_{S}$-integral model $(\mathcal{X},\mathcal{M})=(\mathcal{X},((\mathcal{D}_{1},\dots,\mathcal{D}_{n}),\mathfrak{M}^{\prime}_{\mathcal{C}}))$ of $(X,M)$. This is given by taking the divisors $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ to be the Zariski closures of the divisors $D_{1},\dots,D_{n}$ in $\mathcal{X}_{\mathcal{O}_{\tilde{S}}}$ and by taking $\mathfrak{M}^{\prime}_{\mathcal{C}}$ to be the set obtained by replacing the strata $c$ in $\mathfrak{M}_{\mathcal{C}}$ with their closures $\overline{c}$. (Here we take the Zariski closure of a not necessarily reduced scheme to mean the closure as defined in [Moe24, Proposition 3.6].)
Despite having this canonical choice of an integral model, it is useful to consider other integral models. This discussed in [Moe24, §3.5], of which the main takeaway is as follows. If $f\colon\mathcal{Y}\rightarrow\mathcal{X}$ is a morphism of $\mathcal{O}_{S}$-integral models, then one can pull back $\mathcal{M}$ to get a pair $(\mathcal{Y},f^{-1}\mathcal{M})$ whose points are exactly the preimages of the points on $(\mathcal{X},\mathcal{M})$, but the pair $(\mathcal{Y},f^{-1}\mathcal{M})$ is not necessarily the canonical integral model for $(Y,f^{-1}M)$.
@@
Let $\widetilde{B}\rightarrow B$ and $\hat{B}\rightarrow\widetilde{B}$ be Galois covers of schemes and let $P\colon Y\rightarrow X$ be a morphism of regular schemes over $B$. For any effective divisor $D$ on $X$ and any prime divisor $v$ on $Y$ we have

### v3

@@ 1.4. Evidence for the conjecture
In [32], we prove Conjecture 1.4 for split toric varieties, providing the first results on $\mathcal{M}$-points of bounded height in general.
Theorem 1.12 ([32, Theorem 1.1]).
Let $(X,M)$ be a smooth proper pair over $\mathbb{Q}$, where $X$ is a smooth proper split toric variety and the divisors $D_{1},\dots,D_{n}$ are torus-invariant. Let $\mathcal{X}$ be the $\mathbb{Z}$-integral model of $X$ induced by the fan of $X$ and let $(\mathcal{X},\mathcal{M})$ be the corresponding integral model of $(X,M)$. Then Conjecture 1.4 is true for $(X,M)$ for any big and nef divisor class $L$ on $X$ with toric metrization $\mathcal{L}$ (as defined in [5, Theorem 2.1.6]).
In Section 9, we give further evidence for the conjecture by showing that the conjecture is compatible with the the various results on $\mathcal{M}$-points of bounded height [49, 45, 25, 1]. In Theorem 9.1 we give an asymptotic formula for the number of elements with $m$-full norm in any Galois extension of $\mathbb{Q}$, under assumption of Conjecture 1.4. Furthermore, we show that this asymptotic agrees with the result by Streeter [49, Theorem 1.1], whenever the hypotheses of his result are satisfied. In order to compare our leading constant with the one found by Streeter, we prove Lemma 7.8, giving a description of the algebraic Brauer group of pairs $(X,M)$ over toric varieties in terms of automorphic characters.
@@ 1.5. Structure of the paper
In Section 3, we introduce pairs and $\mathcal{M}$-points, as well as various special cases of these such as Campana points. In Section 4.2, we introduce the Picard group of a pair, as well as its effective cone and its canonical divisor. In the same section, we also prove several results to aid in the computation of the Fujita invariant and the $b$-invariant. In Section 5 we define when a pair is rationally connected, and motivate its role in Conjecture 1.4. Section 6 is devoted to the study of the log-canonical divisor and quasi-Campana pairs, which are pairs loosely resembling the pairs for Campana points. For such pairs, we compute the Fujita invariant and the $b$-invariant, yielding Theorem 1.7 as a special case. In Section 7, we introduce the fundamental group and the Brauer group of a pair, and we study the latter for pairs over toric varieties. In Section 8, we define the leading constant in Conjecture 1.4 using the Brauer group from the previous section. Finally, in Section 9, we prove that the previous results known on $\mathcal{M}$-points of bounded height are compatible with Conjecture 1.4.
@@ Acknowledgments
This article is based on my PhD thesis [31], which I wrote at Utrecht University under the supervision of Marta Pieropan. I would like to thank her for her unwavering support for my work and her various suggestions. I would also like to thank Alec Shute, Sam Streeter, Tim Santens and Damaris Schindler for the useful discussions we had during the workshop on Campana points on toric varieties in Bristol, as well as on other occasions. Finally, I would like to thank Sam Streeter, Tim Santens and Dan Loughran for discussions on the leading constant.
@@ 3. Pairs and $\mathcal{M}$-points
In this section we introduce pairs and the $\mathcal{M}$-points on such pairs. This notion will be based on the notion of pair previously considered in [30, 31]. In this article, we work with a more general notion. The differences are are as follows:
@@
    The parameter set $M$ does not only describe which intersection multiplicities $\mathcal{M}$-points should have, it also can impose to which intersection strata these multiplicities should correspond.
@@
The pairs considered in [30, 31] are the pairs which are split and divisorial, as defined in this section. There are several reasons for considering more general pairs than these. For instance, on a non-split smooth toric variety, the torus-invariant prime divisors are not smooth in general, but their geometric components are smooth. This is significant, as our generalization of Manin’s conjecture requires the divisors $D_{1},\dots,D_{n}$ defining the pair to be smooth.
3.1. Pairs
In this section we will introduce pairs and their $\mathcal{M}$-points. We will introduce these notions in large generality, using the language of scheme theory. While we are mainly interested in $\mathcal{M}$-points over rings of integers, we will also need to consider $\mathcal{M}$-points over $\mathbb{P}^{1}$ in Definition 5.1 to define what it means for a pair to be rationally connected. Furthermore, the general framework presented here has nice functorial properties as seen in Remark 3.19.
In general, we want to consider divisors which need not be not defined over the base, but only over a Galois cover. Let us recall the definition of a Galois cover as in [48, Tag 03SF], generalizing Galois extensions of a field.
@@ 3.1. Pairs
Assumption 3.2.
For the remainder of the section, all schemes will be Noetherian schemes over $B$, and we let $A\rightarrow B$ be a fixed Galois cover with Galois group $G$.
@@
Let $P\colon Y\rightarrow X$ be a morphism of schemes and let $D$ be an effective divisor on $X_{A}$. For any prime (Cartier) divisor $v$ on $Y_{A}$, let $\mathcal{O}_{v}$ be the local ring of $Y_{A}$ at the generic point of $v$, and let $\pi_{v}$ be an uniformizer. Let $P_{v}\colon\operatorname{Spec}\mathcal{O}_{v}\rightarrow X$ be the morphism obtained by composing $P$ with the natural morphism $\operatorname{Spec}\mathcal{O}_{v}\rightarrow Y$. We define the (local) intersection multiplicity of $P$ and $D$ at $v$ to be the extended natural number $n_{v}(D,P)\in\overline{\mathbb{N}}$ such that the fiber product
[…]
is given as $\operatorname{Spec}\mathcal{O}_{v}\times_{X}D=\operatorname{Spec}\mathcal{O}_{v}/(\pi_{v})^{n_{v}(D,P)}$, where we set $(\pi_{v})^{\infty}=0$. Similarly, for an irreducible component $v$ of $Y_{A}$, we set
@@
Note that if the image of $P$ is contained in $D$, then $n_{v}(D,P)=\infty$ for every Cartier divisor $v$ on $Y$. If $D$ is Cartier, the image of $P$ is not in $D$ and $Y$ is integral, then $P^{-1}D$ is the pullback divisor $P^{*}D$ [48, Tag 02OO] and $n_{v}(D,P)$ is the coefficient of $v$ in $P^{*}D$.
@@
Thus $n_{v}(D,P)$ is the local intersection multiplicity of $C$ and $D$ in a point $v$ as defined in [23, Chapter V].
@@
If $X=\mathbb{A}_{\mathbb{Z}}^{2}$, $D_{1}=\{y=0\}$, and $D_{2}=\{y=x^{2}-1\}$, then an integral point $P=(a,b)\in\mathbb{Z}^{2}$ satisfies $n_{p}(D_{1},P)=v_{p}(b)$ and $n_{p}(D_{2},P)=v_{p}(b-a^{2}+1)$ for every prime number $p$, where $v_{p}$ is the $p$-adic valuation. In particular, it follows that $n_{p}(D_{1},P),n_{p}(D_{2},P)>0$ if and only if $b=0\bmod p$ and $a=\pm 1\bmod p$. The intersection multiplicities at $p$ tells us whether the point reduces to $D_{1}\cap D_{2}=\{(1,0),(-1,0)\}$ modulo $p$. However, it does not tell us to which of the two points it reduces to. In order to keep track of this, we introduce the following notation.
@@
We write $\overline{\mathbb{N}}^{n}_{\mathcal{C}}$ for the set of tuples $(\mathbf{m},c)$, where $\mathbf{m}\in\overline{\mathbb{N}}^{n}$ and $c\in\mathcal{C}_{\mathbf{m}}$. The action of $G$ on $\{D_{1},\dots,D_{n}\}$ induces an action on $\overline{\mathbb{N}}^{n}$ by permuting the coordinates, and thus $G$ acts on $\overline{\mathbb{N}}^{n}_{\mathcal{C}}$ by $\sigma(\mathbf{m},c)=(\sigma(\mathbf{m}),\sigma(c))$ for all $\sigma\in G$.
@@
The strata $c$ in Definition 3.7 correspond to the faces of the corresponding Clemens complex (see e.g. [43]) whenever $X$ is a variety and $\sum_{i=1}^{n}D_{i}$ is a normal crossings divisor.
@@
The previous example has a straightforward generalization to toric varieties, see e.g. [30, §5].
@@
Let us compute $\operatorname{mult}^{\mathcal{C}}_{v}(P)$ in Example 3.6 for integral points $P=(a,b)\in\mathbb{Z}^{2}$. For every prime number $p$ we write $\mathbf{m}_{p}=(v_{p}(b),v_{p}(b-a^{2}+1))$, and we will compute the augmented multiplicity of $P$. We consider several cases based on the multiplicity:
  • If $v_{p}(b)=v_{p}(b-a^{2}+1)=0$, then $\operatorname{mult}_{p}^{\mathcal{C}}(a,b)=\{(\mathbf{m}_{p},X)\}$.
  • If $v_{p}(b)>0$ and $v_{p}(b-a^{2}+1)=0$, then $\operatorname{mult}_{p}^{\mathcal{C}}(a,b)=\{(\mathbf{m}_{p},D_{1})\}$,
  • If $v_{p}(b)=0$ and $v_{p}(b-a^{2}+1)>0$, then $\operatorname{mult}_{p}^{\mathcal{C}}(a,b)=\{(\mathbf{m}_{p},D_{2})\}$,
  • If $v_{p}(b),v_{p}(b-a^{2}+1)>0$ and $p$ is odd, then
@@
    where $\overline{a}\in\{1,-1\}$ is the reduction of $a$ modulo $p$.
  • If $v_{p}(b),v_{p}(b-a^{2}+1)>0$ and $p=2$, then
    $\operatorname{mult}_{2}^{\mathcal{C}}(a,b)=\big\{\big(\mathbf{m}_{2},\{(1,0)\}\big),\big(\mathbf{m}_{2},\{(-1,0)\}\big)\big\}.$
@@
If $Y$ is irreducible, then the condition on the irreducible components $v$ of $Y_{A}$ is automatically satisfied if the image of $Y_{A}$ is not contained in any of the divisors.
@@
The assumption that $Y$ is regular and Noetherian implies that every divisor on $Y$ is a sum of prime Cartier divisors by [48, Tag 0BCP]. Therefore, the values $\operatorname{mult}_{v}(P)$ for all $v$ as above together determine the inverse images $P^{-1}D_{1},\dots,P^{-1}D_{n}\subset Y_{A}$.
For many pairs considered in this article, no condition is imposed on the strata $c$ but only on the intersection multiplicity. To make the notation less cumbersome in this case, we introduce the following definition.
@@
The term “divisorial” is used, because in this case the criterion for a morphism $P\colon Y\rightarrow X$ to be an $M$-point depends solely on the intersection multiplicities of $P$ with the divisors $D_{1},\dots,D_{n}$, and not on their intersection strata.
@@
The definition of a divisorial split pair given here strongly resembles the definition of a pair previously used by the author in [30, Definition 3.1], but there are two minor differences. In this article, we assume for simplicity that $D_{1},\dots,D_{n}$ are divisors rather than arbitrary closed subschemes. Furthermore, in the other article, we imposed a weak condition on the set $\mathfrak{M}$ ensuring that the $\mathcal{M}$-points over $\mathbb{Z}$ are a subset of the $\mathcal{M}$-points over $\mathbb{Q}$. This condition is not necessary in the framework used in this paper, as the $\mathcal{M}$-point condition used here also takes generic points into account.
@@
If the divisors $D_{1},\dots,D_{n}$ are chosen such that $D_{1}+\dots+D_{n}$ is a strict normal crossings divisor as in [48, Tag 0BI9], then [48, Tag 0BIA] implies that the connected components of the intersections $\bigcap_{i\in I}D_{i}$ are all irreducible for all nonempty $I\subset\{1,\dots,n\}$. Thus it follows that for any scheme $Y$ over $B$ with a prime divisor (or connected component) $v$, the function $\operatorname{mult}_{v}^{\mathcal{C}}$ can be regarded as a function $X(Y)\rightarrow\overline{\mathbb{N}}^{n}_{\mathcal{C}}$. In particular, the condition at $v$ for $P\colon Y\rightarrow X$ to be an $M$-point simplifies to $\operatorname{mult}^{\mathcal{C}}_{v}(P)\in\mathfrak{M}_{\mathcal{C}}$.
@@
for $a_{v}\in\overline{\mathbb{N}}$. Here $a_{v}$ is the multiplicity of $v^{\prime}$ in the pullback of $v$ to $Y^{\prime}_{A}$ if $f(Y^{\prime})\not\subset v$ and $a_{v}=\infty$ otherwise. This can be seen as follows: if the image of $P\circ f$ is not contained in an effective Cartier divisor $D\subset X$, then the scheme-theoretic fiber $(P\circ f)^{-1}D=(P\circ f)^{*}D$ is a Cartier divisor on $Y^{\prime}$. The desired identity now follows from $(P\circ f)^{*}D=f^{*}P^{*}D$ together with the linearity of $f^{*}$.
@@ 3.2. Pairs over number fields
Let $(X,M)$ be a pair over a number field $K$, where $X$ is a proper variety. For an integral model $\mathcal{X}$ of $X$ over $\mathcal{O}_{S}$, where $S$ is a finite set of places of $K$, there is a natural choice of an $\mathcal{O}_{S}$-integral model $(\mathcal{X},\mathcal{M})=(\mathcal{X},((\mathcal{D}_{1},\dots,\mathcal{D}_{n}),\mathfrak{M}^{\prime}_{\mathcal{C}}))$ of $(X,M)$. This is given by taking the divisors $\mathcal{D}_{1},\dots,\mathcal{D}_{n}$ to be the Zariski closures of the divisors $D_{1},\dots,D_{n}$ in $\mathcal{X}_{\mathcal{O}_{\tilde{S}}}$ and by taking $\mathfrak{M}^{\prime}_{\mathcal{C}}$ to be the set obtained by replacing the strata $c$ in $\mathfrak{M}_{\mathcal{C}}$ with their closures $\overline{c}$. (Here we take the Zariski closure of a not necessarily reduced scheme to mean the closure as defined in [30, Proposition 3.6].)
Despite having this canonical choice of an integral model, it is useful to consider other integral models. This discussed in [30, §3.5], of which the main takeaway is as follows. If $f\colon\mathcal{Y}\rightarrow\mathcal{X}$ is a morphism of $\mathcal{O}_{S}$-integral models, then one can pull back $\mathcal{M}$ to get a pair $(\mathcal{Y},f^{-1}\mathcal{M})$ whose points are exactly the preimages of the points on $(\mathcal{X},\mathcal{M})$, but the pair $(\mathcal{Y},f^{-1}\mathcal{M})$ is not necessarily the canonical integral model for $(Y,f^{-1}M)$.
@@
Let $\widetilde{B}\rightarrow B$ be a Galois cover of schemes and let $P\colon Y\rightarrow X$ be a morphism of regular schemes over $B$. For any effective divisor $D$ on $X$ and any prime divisor $v$ on $Y$ we have